ServiceAgent1:

The teacher bought new crayons for her class. She bought 7 packs of 9 crayons. She already has 17 crayons in her classroom. She wants to divide the crayons evenly between her 10 students. Write an equation using (c) as the variable and show your work to determine how many crayons each student will receive.

carmelle:

First we start by doing \[7 \times 9 = ?\] what does that equal?

ServiceAgent1:

63

carmelle:

correct, now we do \[63 + 17 = ?\] What does that equal?

carmelle:

@serviceagent1

ServiceAgent1:

80

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @carmelle correct, now we divide it by 10 so..\[80\div10=c\] \(\color{#0cbb34}{\text{End of Quote}}\)

ServiceAgent1:

8

carmelle:

correct, so each student gets 8 crayons the final equation would be \[\frac{ (7\times9) + 17 }{ 10 } = c\]

ServiceAgent1:

okt hank u

carmelle:

np